OpenStudy (anonymous):

MEDAL & FAN 5. Using the points (3, 4) and (-5, 8) find the following: a. Slope. b. Point-Slope form of the equation. c. Slope-intercept form of the equation

OpenStudy (anonymous):

I think the slope is 1/-2 but i could be wrong... I'm not too good in this aspect of math.

OpenStudy (anonymous):

yep u r right

OpenStudy (anonymous):

okay but what about the rest of the problem?

OpenStudy (anonymous):

i think the point slope form means that its probably a negative slope

OpenStudy (anonymous):

and i think you should try drawing it it might help

OpenStudy (campbell_st):

ok... the slope is correct...-1/2 the point slope form is \[y - y_{1} = m(x - x_{1})\] where \[(x_{1}, y_{1}) \] is the given point so you just substitute the given information.

OpenStudy (stamp):

y - y1 = m (x - x1) (x1, y1) being an order pair on the line You have identified the slope m to be equal to -1/2

OpenStudy (campbell_st):

slope intercept form is when you use the point slope form and then distribute the -1/2 and collect like terms and get in in the form \[y = mx + b\] hope it helps

OpenStudy (stamp):

^ What @campbell_st provided is on point

OpenStudy (anonymous):

Thanks guys! Really helped (:

OpenStudy (campbell_st):

just choose 1 of the 2 points given, for the point slope form. It doesn't matter which one as they both will give the same answer

OpenStudy (anonymous):

How do I get Slope Intercept?

OpenStudy (campbell_st):

well after you get the point slope form... just distribute and then make y the subject e.g. if I have a slope of 5 and point (1, 2) the point slope is y - 2 = 5(x - 1) then distribute y - 2 = 5x - 5 now make y the subject by adding 2 to both sides of the equation y = 5x - 3 this is slope intercept form

OpenStudy (anonymous):

I'm still lost... :/ Sorry

OpenStudy (stamp):

For future reference, to recall what slope-intercept form is and looks like, look at the name of the form Slope - intercept OR y = mx + b m is your slope, and recall that b represents the intercept. So y = mx + b, it tells you your slope and your intercept. Hence the name, slope-intercept form.

OpenStudy (anonymous):

So in this case it'd be: y= -1/2 x + 4 ?

OpenStudy (campbell_st):

which point are you using (3, 4) or (-5, 8)

OpenStudy (anonymous):

(3,4)

OpenStudy (campbell_st):

ok... so the point slope form is y - 4 = -1/2(x - 3) does that make sense...?

OpenStudy (campbell_st):

ok... so when you distribute the -1/2 its y - 4 = -1/2 x + 3/2 is that ok...?

OpenStudy (anonymous):

I got that on my own. but Slope Intercept form is what I'm trying to find.

OpenStudy (stamp):

Knowing the slope m is -1/2, you can go from writing y = mx + b to the incomplete equation y = -1/2 x + b Using the ordered pair provided (3, 4), you can substitute the x and y values into the equation to help find b, so 4 = -1/2 (3) + b 4 = -3/2 + b 8/2 + 3/2 = b b = 11/2 So, knowing b, you can take your incomplete equation y = -1/2 x + b and declare y = -1/2 x + 11/2

OpenStudy (anonymous):

you guys have 2 different answers :/

OpenStudy (campbell_st):

no we don't if you have y - 4 = -1/2 x + 3/2 now add 4 to both sides of the equation. then you will have the slope intercept form

OpenStudy (stamp):

If our equation y = -1/2 x + 11/2 is correct, the ordered pair provided (-5, 8) should satisfy the equation. Let us try it: 8 = -1/2 (-5) + 11/2 8 = 5/2 + 11/2 8 = 16/2 8 = 8 The equation is verified.

OpenStudy (anonymous):

OH! Okay. Thanks guys!